PeppaPig:

Camille's Seed Haven claims that 90% of its sunflower seeds will germinate. Suppose the company's claim is true. Seth buys a packet with 30 sunflower seeds from Camille's Seed Haven and plants them in his garden. What is the probability that exactly 23 seeds will germinate?

jsvsllys:

I can help I did this before

PeppaPig:

thx can u explain

jsvsllys:

sure

jsvsllys:

This is a binomial probability distribution function because it's asking for an exact value. You won't use a Animeulative function. The formula for a bionomial probability distribution is nCr*p^r*q^(n-r) such that n=total trials, r=successes desired, p=probability of a success, q=probability of failure or 1-p. in this scenario n=30, r=23, p=0.9, q=1-0.9=0.1 Plug in the values, compute it, and you'll get 0.0180. That's the probability 23 seeds will germinate.

PeppaPig:

thx

PeppaPig:

r u sure this is my last chance

jsvsllys:

yeah, I did this before

PeppaPig:

k